Darfangel:

I need help pwease! :0

Darfangel:

Fill in the blank question. If f(x)=3x+2 and g(x)=x2−x, find the value.

Darfangel:

If you smart explain to me the answer sailor you pretty smart i know!

Sock:

yes if anyone can answer this it would be sailor

Sock:

now if you'll excuse me im off the hunt down a person with the same name as me

NoodlesAndRiceYT:

@sock wrote:
yes if anyone can answer this it would be sailor
no presser sailor

Darfangel:

1 attachment
Darfangel:

thats the problem!

NoodlesAndRiceYT:

"how to screenshot on windows" i cant- 💀😂

Sailor:

Ok, so because you have to find F(-2), you have to plug it into F(x)=3x+2.

Darfangel:

okie

Sailor:

So your equation would be F(-2)=3(-2)+2

Darfangel:

would it be -4

Sailor:

Yeah.

Sailor:

Then for g(x)=x2−x, is the x2 squared?

Darfangel:

okie thx i was stuck on that question for a while!

Darfangel:

@sailor

Sailor:

Ok

Sailor:

If you need help with anything else, let me know.